(e^2x + 4) dy/dx = y


1
Expert's answer
2021-09-10T00:09:47-0400

"\frac{dx}{e^{2x}+4}=\frac{dy}{y}"

"Let z=e^{2x}+4"

"dz=2e^{2x}dx=2(z-4)dx"

"dx=\frac{dz}{2(z-4)}"

"\frac{dy}{y}=\frac{dx}{z}=\frac{dz}{2z(z-4)}=1/8(\frac{1}{z-4}-\frac{1}{z})dz"

"lny=1\/8(ln(z-4)-lnz)+C"

"lny=1\/8(2x-ln(e^{2x}+4))+C"